# First-Order Sparse Convex Optimization: Better Rates with Sparse Updates

### v2

@@ Abstract
It was recently established that for convex optimization problems with a sparse optimal solution (may it be entry-wise sparsity or matrix rank-wise sparsity) it is possible to have linear convergence rates which depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipchitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of the optimal solution. However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving also the runtime of each iteration, which may still be prohibitively high for high-dimensional problems. In this work, we establish that linear convergence rates which depend on this improved condition number can be obtained using only sparse updates, which may result in overall significantly improved running times. Moreover, our methods are considerably easier to implement.
@@ 1 Introduction
[…]
Example 2: trace-constrained minimization for low-rank matrix recovery.
$\min_{{\mathbf{X}}\in\mathbb{S}^{n}:~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R}f({\mathbf{X}}),$ (2)
where $\mathbb{S}^{n}$ is the space of real symmetric $n\times{}n$ matrices and the relation ${\mathbf{X}}\succeq 0$ denotes ${\mathbf{X}}$ is positive semidefinite. Here we also assume $f$ is convex and continuously differentiable, and that Problem (2) indeed admits a unique optimal solution ${\mathbf{X}}^{*}$ such that $\textrm{rank}({\mathbf{X}}^{*})<<n$.
[…]
In recent works [16, 14], it was established that for convex relaxations of sparse optimization problems such as Problems (1) and (2), and under the quite plausible assumption that the optimal solution is indeed sparse/low-rank, it is possible to obtain linear convergence rates which depend on a mixed-norm condition number of the form $\frac{\beta_{1}s}{\alpha_{2}}$, where $\beta_{1}$ is the Lipchitz continuity constant of the gradient w.r.t. the $\ell_{1}$ norm, $\alpha_{2}$ is the quadratic growth constant w.r.t. the $\ell_{2}$ norm, and $s$ is an upper-bound on the sparsity of the optimal solution. This condition number could indeed be far better than the pure $\ell_{2}$ or pure $\ell_{1}$ condition numbers. For instance, in the context of Problem (1), consider the case that $f$ is convex quadratic of the following very simple form:
@@
Indeed if $\|{{\mathbf{x}}^{*}}\|_{1}\leq R$, ${\mathbf{x}}^{*}$ is the unique optimal solution to Problem (1) for this $f$. One can easily observe that the Euclidean condition number for this problem is given by $\frac{\beta_{2}}{\alpha_{2}}=\frac{\lambda_{\max}({\mathbf{A}})}{\lambda_{\min}({\mathbf{A}})}=\frac{n+\alpha}{\alpha}$. On the other-hand, a simple calculation shows that the $\ell_{1}$-Lipchitz continuity constant of the gradient is given by $\beta_{1}=\max_{i,j}|{{\mathbf{A}}_{i,j}}|=1+\alpha$. To bound the $\ell_{1}$-quadratic growth constant we observe that due the the structure of the matrix ${\mathbf{A}}$,
@@
In particular, assuming ${\mathbf{x}}^{*}\geq 0$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$ (i.e., ${\mathbf{x}}^{*}$ lies on the boundary of the $\ell_{1}$ ball, as often expected) we have that
$\forall{\mathbf{x}}\in\{{\mathbf{y}}\in\mathbb{R}^{n}~|~{\mathbf{y}}\geq 0,\|{{\mathbf{y}}}\|_{1}=R\}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}.$
@@
Thus, in worst-case, the $\ell_{1}$-quadratic growth constant $\alpha_{1}$ is not better than $\frac{\alpha}{n}$, which implies that the $\ell_{1}$-condition number is $\frac{\beta_{1}}{\alpha_{1}}=\frac{n+n\alpha}{\alpha}$, which is no better than the Euclidean one.
@@
However, in the example above, if ${\mathbf{x}}^{*}$ satisfies $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$, the sparsity-based mixed-norm condition number satisfies $\frac{\beta_{1}s}{\alpha_{2}}=\frac{s(1+\alpha)}{\alpha}$, which is in particular independent of the ambient dimension, and thus may lead to far-better (worst-case) convergence rates for suitable first-order methods.
@@
The example in Eq. (3) could be directly extended also to Problem (8) by taking $f({\mathbf{X}})=\frac{\alpha}{2}\|{{\mathbf{X}}-{\mathbf{X}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\textrm{Tr}({\mathbf{X}}-{\mathbf{X}}^{*})}\right)^{2}$, where $\|{\cdot}\|_{2}$ here is the Euclidean norm in $\mathbb{S}^{n}$, and instead of the entry-wise $\ell_{1}$ norm we consider the Schatten 1-norm (i.e., the $\ell_{1}$ norm w.r.t. the eigenvalues), and $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$.
The recent works [16, 14]111while [16, 14] consider a stochastic convex optimization setting in which the objective function $f$ is given as an expectation w.r.t. unknown distribution from which we can sample, here our focus is on the standard deterministic first-order convex optimization model indeed lead to convergence rates of the form $\tilde{O}\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ [16] and even an accelerated rate of the form $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$ [14] (the $\tilde{O}(\cdot)$ notation hides certain poly-logarithmic factors), by applying a restarted non-Euclidean gradient method222these methods rely on Bregman proximal gradient computations with the Bregman distance induced by a squared $\ell_{p}$ norm for an appropriate choice of $p$ that best leverages the strong convexity w.r.t. the $\ell_{1}$ norm or its accelerated variant, together with hard-thresholding steps to sparsify intermediate iterates, see for instance Corollary 5 in [14].
@@
While these non-Euclidean gradient methods lead to superior convergence rates, they do not leverage the sparsity of optimal solutions towards additional computational gains, and in particular may still require prohibitive running times per iteration when the dimension is large. For instance, in the case of Problem (1), these may suffer from prohibitive running time for updating the gradient direction from one iteration to the next. In the case of Problem (2) the time required by these methods to project onto the positive semidefinite cone ($O(n^{3})$ in standard implementations) may be prohibitive.
@@
In this work we present a new and simple approach towards obtaining linear convergence rates that depend on the improved condition number $\frac{\beta_{1}s}{\alpha_{2}}$ (though our method only obtains rates that are linear in it and not accelerated rates which depend only on its square root). The main novel aspect is that our algorithm only applies sparse updates (with the same level of sparsity assumed for the optimal solution), which often allows for significant improvements in the runtime of each iteration. In particular, our algorithm updates its iterates similarly to the classical Frank-Wolfe algorithm [15]:
${\mathbf{x}}_{t+1}\leftarrow(1-\eta){\mathbf{x}}_{t}+\eta{\mathbf{v}}_{t}\quad\textrm{for some}\quad\eta\in[0,1],$ (4)
@@
where ${\mathbf{v}}_{t}$ is a feasible point with the same level of sparsity assumed for ${\mathbf{x}}^{*}$.
@@
For instance, in case of Problem (1) with a quadratic objective function $f({\mathbf{x}})={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, the fact that ${\mathbf{v}}_{t}$ is $s$-sparse, reduces the time to update the gradient direction from $O(n^{2})$ to $O(sn)$. In the case of Problem (2), since $\textrm{rank}({\mathbf{v}}_{t})\leq s$, computing it won’t require a full eigen-decomposition ($O(n^{3})$ time in standard implementations), but only computing the $s$ leading components in the eigen-decomposition, which is far more efficient.
@@
Two additional advantages which make our approach considerably simpler (compared to [16, 14]) are: I. it does not require restarts, i.e., it is a single-loop algorithm, and II. the updates in our algorithm (the way ${\mathbf{v}}_{t}$ in Eq. (4) is computed) are Euclidean, which often leads to significantly simpler computations.
@@
Compared to [16, 14], we do have an additional mild assumption, which in the case of Problems (1) and (2) above, requires that the optimal solution is on the boundary in the sense that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ or $\textrm{Tr}({\mathbf{X}}^{*})=R$, respectively. We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm regularization is redundant).
@@
The rest of the paper is organized as follows. In Section 2 we present a unified setup and optimization problem which captures the spare convex optimization problems we are interested in, a generic algorithm for solving this problem, and its convergence analysis. In Section 3 we instantiate our generic algorithm in the special cases of Problems (1) and (2). In Section 4 we give a simple numerical demonstration for the performance of our method on Problem (1). Finally, in Section 5 we outline several open questions.
@@ 2 Unified Setup, Algorithm, and Convergence Analysis
[…]
@@
We let $\|{{\mathbf{x}}}\|_{2}=\sqrt{\langle{{\mathbf{x}},{\mathbf{x}}}\rangle}$ denote the Euclidean norm.
We shall assume Problem (5) admits a unique optimal solution ${\mathbf{x}}^{*}$, and satisfies a quadratic growth condition with parameter $\alpha_{2}>0$ w.r.t. the Euclidean norm, i.e.,
$\forall{\mathbf{x}}\in\mathcal{K}:\quad\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}\leq\frac{2}{\alpha_{2}}\left({f({\mathbf{x}})-f({\mathbf{x}}^{*})}\right).$ (6)
We now make an additional assumption that further specializes Problem (5). In Section 3 we show that in the case of Problems (1) and (2) (and certain generalizations), Assumption 1 is satisfied once we make the plausible assumption that the norm constraint is tight for the optimal solution ${\mathbf{x}}^{*}$.
@@
There exists a closed subset $\mathbb{E}_{s}\subset\mathbb{E}$ such ${\mathbf{x}}^{*}\in\mathcal{K}_{s}:=\mathcal{K}\cap\mathbb{E}_{s}$, and there exist scalars $s,s_{\mathcal{K}}^{*}\geq 1$ such that,
@@
    for all ${\mathbf{x}},{\mathbf{y}}\in\mathbb{E}_{s}$: $\|{{\mathbf{x}}-{\mathbf{y}}}\|\leq\sqrt{s}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}$,
@@
    for all ${\mathbf{x}}\in\mathcal{K}$: $\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|\leq\sqrt{s^{*}_{\mathcal{K}}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}$,
@@
    for all ${\mathbf{x}}\in\mathcal{K}$: $\arg\min\nolimits_{{\mathbf{y}}\in\mathbb{E}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|=\arg\min\nolimits_{{\mathbf{y}}\in\mathcal{K}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|_{2}$ (note the argmin need not be a singleton),
In the context of Problems (1) and (2) we should think of $\mathbb{E}_{s}$ as the restriction of the space $\mathbb{E}$ to vectors which are $s$-sparse (in some sense), and accordingly $\mathcal{K}_{s}$ is the restriction of $\mathcal{K}$ to $s$-sparse vectors. Additionally, while not part of Assumption 1, we shall implicitly assume that it is computationally efficient to compute a Euclidean projection onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$ (in case the projection is not a singleton, compute one element in the set of possible projections). Accordingly, we denote by $\Pi_{\mathbb{E}_{s}}[\cdot]$ and $\Pi_{\mathcal{K}_{s}}[\cdot]$ the projection operation onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$, respectively.
We can now present our algorithm for solving the generic Problem (5) (under the assumptions listed above). Our algorithm is composed of three simple ingredients. First, we use a Frank-Wolfe like scheme [15] in which the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ are updated by taking a convex combination between the current feasible iterate and a new feasible point, as already shown in Eq. (4). Second, and differently from the classical Frank-Wolfe method in which the new point ${\mathbf{v}}_{t}$ is the output of a linear optimization oracle, here we take it to be a Euclidean proximal gradient step w.r.t. to restricted set $\mathcal{K}_{s}$ (instead of w.r.t. the original set $\mathcal{K}$). This is similar to approach taken in previous works [1, 10, 12, 13] (also referred to as a weak proximal oracle [10, 13]). Third, the $\ell_{2}$-term in this proximal gradient step is not w.r.t. to the current iterate ${\mathbf{x}}_{t}$, but with respect to its projection onto $\mathbb{E}_{s}$, which should be understood as a hard-thresholding (or sparsification) of ${\mathbf{x}}_{t}$.
@@
  ${\mathbf{v}}_{t}\leftarrow\arg\min\nolimits_{{\mathbf{v}}\in\mathcal{K}_{s}}\langle{{\mathbf{v}}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+s\beta\eta\|{{\mathbf{v}}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}$equivalently, ${\mathbf{v}}_{t}\leftarrow\Pi_{\mathcal{K}_{s}}[\widehat{{\mathbf{x}}}_{t}-\frac{1}{2s\beta\eta}\nabla{}f({\mathbf{x}}_{t})]$}
  ${\mathbf{x}}_{t+1}\leftarrow(1-\eta){\mathbf{x}}_{t}+\eta{\mathbf{v}}_{t}$

### v3

@@ Abstract
It was recently established that for convex optimization problems with sparse optimal solutions (be it entry-wise sparsity or matrix rank-wise sparsity) it is possible to design first-order methods with linear convergence rates that depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipschitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of optimal solutions. However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving the runtime of each iteration as well, which may still be prohibitively high for high-dimensional problems. In this work, we establish that linear convergence rates which depend on this improved condition number can be obtained using only sparse updates, which may result in overall significantly improved running times. Moreover, our methods are considerably easier to implement.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is convex and continuously differentiable, and we assume that all optimal solutions have sparsity (number of non-zero entries) at most $s<<n$.
Example 2: Nuclear norm-constrained minimization for low-rank matrix recovery.
$\min_{{\mathbf{X}}\in\mathbb{R}^{m\times n}:~\|{{\mathbf{X}}}\|_{1}\leq R}f({\mathbf{X}}),$ (2)
where here we let $\|{\cdot}\|_{1}$ denote the matrix nuclear norm in $\mathbb{R}^{m\times n}$, i.e., sum of singular values.
Here we also assume $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ is convex and continuously differentiable, and that all optimal solutions have rank at most $s<<\min\{m,n\}$.
With regard to both problems, let us denote by $\beta$ the Lipschitz continuity constant of the gradient w.r.t. some norm $\|{\cdot}\|$, i.e., for all ${\mathbf{x}},{\mathbf{y}}$ we have that $\|{\nabla{}f({\mathbf{x}})-\nabla{}f({\mathbf{y}})}\|_{*}\leq\beta\|{{\mathbf{x}}-{\mathbf{y}}}\|$, where $\|{\cdot}\|_{*}$ denotes the norm dual to $\|{\cdot}\|$. Suppose also that $f$ satisfies a lower curvature condition such as strong convexity, or the weaker quadratic growth condition [22], with constant $\alpha>0$ w.r.t. the same norm, i.e., that for any feasible $\mathbf{x}$} it holds that $\min_{{\mathbf{x}}^{*}\in\mathcal{X}^{*}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|^{2}\leq\frac{2}{\alpha}\left({f({\mathbf{x}})-f({\mathbf{x}}^{*})}\right)$, where we denote by $\mathcal{X}^{*}$ the set of all optimal solutions. In this setup, known gradient methods can produce an $\epsilon$-approximated solution (in function value) in worst-case number of iterations that scales like $\frac{\beta}{\alpha}\log{1/\epsilon}$, or even $\sqrt{\frac{\beta}{\alpha}}\log{1/\epsilon}$ when accelerated methods are used [24, 2, 22]. The ratio $\beta/\alpha$ is the condition number of $f$ w.r.t. the norm $\|{\cdot}\|$ and is naturally dependent on the choice of norm.
In recent works [17, 15], it was established that for convex relaxations of sparse optimization problems such as Problems (1) and (2), and under the quite plausible assumption that all optimal solutions are indeed sparse/low-rank, it is possible to design gradient methods that obtain linear convergence rates which depend on a mixed-norm condition number of the form $\frac{\beta_{1}s}{\alpha_{2}}$, where $\beta_{1}$ is the Lipschitz continuity constant of the gradient w.r.t. the $\ell_{1}$ norm, $\alpha_{2}$ is the quadratic growth constant w.r.t. the $\ell_{2}$ norm, and $s$ is an upper-bound on the sparsity of optimal solutions. This condition number could indeed be far better than the pure $\ell_{2}$ or pure $\ell_{1}$ condition numbers. For instance, in the context of Problem (1), consider the case that $f$ is convex quadratic of the following very simple form:
@@
Indeed if $\|{{\mathbf{x}}^{*}}\|_{1}\leq R$, $\mathbf{x}$$^{*}$ is the unique optimal solution to Problem (1) for this $f$. One can easily observe that the Euclidean condition number for this problem is given by $\frac{\beta_{2}}{\alpha_{2}}=\frac{\lambda_{\max}({\mathbf{A}})}{\lambda_{\min}({\mathbf{A}})}=\frac{n+\alpha}{\alpha}$. On the other hand, a simple calculation shows that the $\ell_{1}$-Lipschitz continuity constant of the gradient is given by $\beta_{1}=\max_{i,j}|{{\mathbf{A}}_{i,j}}|=1+\alpha$. To bound the $\ell_{1}$-quadratic growth constant we observe that due to the structure of the matrix $\mathbf{A}$},
@@
In particular, assuming ${\mathbf{x}}^{*}\geq 0$ we have that
$\forall{\mathbf{x}}\in\{{\mathbf{y}}\in\mathbb{R}^{n}~|~{\mathbf{y}}\geq 0,\|{{\mathbf{y}}}\|_{1}=\|{{\mathbf{x}}^{*}}\|_{1}\}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}.$
@@
This implies that the $\ell_{1}$-quadratic growth constant $\alpha_{1}$ is not better than $\frac{\alpha}{n}$, which implies that the $\ell_{1}$-condition number is $\frac{\beta_{1}}{\alpha_{1}}=\frac{n+n\alpha}{\alpha}$, which is no better than the Euclidean one.
@@
However, in the example above, if $\mathbf{x}$$^{*}$ satisfies $\mathrm{nnz}({\mathbf{x}}^{*})\leq s<<n$, the sparsity-based mixed-norm condition number satisfies $\frac{\beta_{1}s}{\alpha_{2}}=\frac{s(1+\alpha)}{\alpha}$, which is in particular independent of the ambient dimension, and thus may lead to far-better (worst-case) convergence rates for suitable first-order methods.
@@
The example in Eq. (3) could be directly extended also to Problem (2) by considering the space $\mathbb{R}^{n\times n}$ and taking $f({\mathbf{X}})=\frac{\alpha}{2}\|{{\mathbf{X}}-{\mathbf{X}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\mathrm{Tr}({\mathbf{X}}-{\mathbf{X}}^{*})}\right)^{2}$ with $\mathbf{X}$$^{*}$ satisfying $\mathrm{rank}({\mathbf{X}}^{*})\leq s<<n$, and letting $\|{\cdot}\|_{2}$ here denote the Euclidean (Frobenius) norm, and instead of the entry-wise $\ell_{1}$ norm, we consider the matrix nuclear norm ($\ell_{1}$ norm w.r.t. the vector of singular values).
[…]
@@
While these non-Euclidean gradient methods lead to superior convergence rates, they do not leverage the sparsity of optimal solutions towards additional computational gains, and in particular may still require prohibitive running times per iteration when the dimension is very high. In the case of Problem (1), these may suffer from prohibitive running time for updating the gradient direction from one iteration to the next. In the case of Problem (2), the time required by these methods to project onto the nuclear norm ball ($O(\min\{m,n\}^{2}\max\{m,n\})$ time, in standard implementations, to compute a full compact-form SVD) may be prohibitive.
@@
In this work we present a new and simple approach towards obtaining linear convergence rates that depend on the mixed-norm condition number $\frac{\beta_{1}s}{\alpha_{2}}$ (though our method only obtains rates that are linear in it and not accelerated rates that scale with its square root). The main novel aspect is that our algorithm only applies sparse updates (with the same level of sparsity assumed for the optimal solutions), which often allows for significant improvements in the runtime of each iteration. In particular, our algorithm updates its iterates similarly to the classical Frank-Wolfe algorithm [16]:
${\mathbf{x}}_{t+1}\leftarrow(1-\gamma){\mathbf{x}}_{t}+\gamma{\mathbf{v}}_{t}\quad\textrm{for some}\quad\gamma\in[0,1],$ (4)
@@
where ${\mathbf{v}}_{t}$ is a feasible point with the same level of sparsity assumed for the optimal solutions.
@@
For instance, in case of Problem (1) with a quadratic objective function $f({\mathbf{x}})={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, the fact that ${\mathbf{v}}_{t}$ is $s$-sparse, reduces the time to update the gradient direction from $O(n^{2})$ to $O(sn)$. In the case of Problem (2), the fact that $\mathrm{rank}({\mathbf{v}}_{t})\leq s$ implies that computing it will not require a full compact-form SVD ($O(\min\{m,n\}^{2}\max\{m,n\})$ time in standard implementations), but only computing the $s$ leading components in the SVD, which is far more efficient.
@@
Two additional advantages which make our approach considerably simpler (compared to [17, 15]) are: I. it does not require restarts, i.e., it is a single-loop algorithm, and II. the updates in our algorithm (the way ${\mathbf{v}}_{t}$ in Eq. (4) is computed) are Euclidean, which often leads to significantly simpler implementations.
@@
Compared to [17, 15], we do have an additional mild assumption, which in the case of Problems (1) and (2) requires that all optimal solutions are on the boundary, i.e., that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ for any optimal solution $\mathbf{x}$$^{*}$. We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm constraint is redundant).
Aside from our original method we also make the observation that for Problem (1), another method that is able to attain linear convergence with the mixed-norm condition number $\frac{\beta_{1}s}{\alpha_{2}}$ that uses only sparse updates is the well-known away-steps Frank-Wolfe (AFW) method [14, 20]. In Section A we show that, assuming all optimal solutions to Problem (1) are $s$-sparse and lie on the boundary (same assumption made for our proposed method), AFW has linear convergence of the form $O\left({\frac{\beta_{1}sR^{2}}{\alpha_{2}}\log{1/\epsilon}}\right)$. Note the additional dependence on the squared radius $R^{2}$. In particular, under the plausible assumption that $R=\Omega(s)$ (e.g., entries in optimal solutions are $\Theta(1)$), this rate becomes $\Omega\left({\frac{\beta_{1}s^{3}}{\alpha_{2}}\log{1/\epsilon}}\right)$. Note also that AFW does not extend to handle Problem (2).
Algorithm required parameters #iterations to $\epsilon$ error iteration time
minimization over $\ell_{1}$-ball with $f({\mathbf{x}})=\frac{1}{2}{\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, ${\mathbf{A}}\succeq 0$, $\mathrm{nnz}({\mathbf{x}}^{*})\leq s~\forall{\mathbf{x}}^{*}\in\mathcal{X}^{*}$
Away-Step Frank-Wolfe with line-search [14, 20] (see Section A) - $\frac{\beta_{1}sR^{2}}{\alpha_{2}}$ $O(n)$
Restarted Acc. Grad. with hard-thresholding [15] (see Section B) $s,\beta_{1},\alpha_{2}$ $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ $O(n^{2})$
Our Algorithm 1 (see Corollary 1) $s,\alpha_{2}$ $\frac{\beta_{1}s}{\alpha_{2}}$ $O(sn)$
minimization over matrix nuclear norm ball, $\mathrm{rank}({\mathbf{X}}^{*})\leq s~\forall{\mathbf{X}}^{*}\in\mathcal{X}^{*}$
Restarted Acc. Grad. with hard-thresholding [15] (see Section B) $s,\beta_{1},\alpha_{2}$ $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ full SVD of $m\times n$ matrix
Our Algorithm 1 (see Corollary 2) $s,\alpha_{2}$ $\frac{\beta_{1}s}{\alpha_{2}}$ rank-$s$ SVD of $m\times n$ matrix
Table 1: Comparison of first-order methods with linear convergence rates that scale with the mixed-norm condition number $\frac{\beta_{1}s}{\alpha_{2}}$ for Problems (1) and (2), omitting logarithmic factors and universal constants.
From Table 1 we can see for instance that in case of Problem (1) with a quadratic objective, whenever $s<<\min\{R^{2},n^{2/3}\left({\alpha_{2}/\beta_{1}}\right)^{1/3}\}$, the overall runtime of our method is the state-of-the-art.
@@
The rest of the paper is organized as follows. In Section 2 we present a unified setup and optimization problem which captures the sparse convex optimization problems we are interested in, a generic algorithm for solving this problem, and its convergence analysis. In Section 3 we instantiate our generic algorithm in the special cases of Problems (1) and (2). In Section 4 we give a simple numerical demonstration for the performance of our method on Problem (1). Finally, in Section 5 we outline several open questions.
@@ 2 Unified Setup, Algorithm, and Convergence Analysis
where $\mathcal{K}$ is a convex and compact subset of a Euclidean vector space $\mathbb{E}$ equipped with an inner-product $\langle{\cdot,\cdot}\rangle$, $f$ is convex over $\mathcal{K}$ with Lipschitz continuous gradient with parameter $\beta$ w.r.t. some norm $\|{\cdot}\|$, i.e.,
@@
We let $f^{*}$ denote the optimal value of (5) and we let $\mathcal{X}^{*}\subseteq\mathcal{K}$ denote the set of minimizers. We let $\|{{\mathbf{x}}}\|_{2}=\sqrt{\langle{{\mathbf{x}},{\mathbf{x}}}\rangle}$ denote the Euclidean norm.
We shall assume Problem (5) satisfies a quadratic growth condition with parameter $\alpha_{2}>0$ w.r.t. the Euclidean norm, i.e.,
$\forall{\mathbf{x}}\in\mathcal{K}:\quad\min_{{\mathbf{x}}^{*}\in\mathcal{X}^{*}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}\leq\frac{2}{\alpha_{2}}\left({f({\mathbf{x}})-f({\mathbf{x}}^{*})}\right).$ (6)
We now make an additional assumption that further specializes Problem (5). In Section 3 we show that in the case of Problems (1) and (2) (and certain generalizations), Assumption 1 is satisfied once we make the plausible assumption that the norm constraint is tight for all optimal solutions.
@@
There exists a closed subset $\mathbb{E}_{s}\subset\mathbb{E}$ such that $\mathcal{X}^{*}\subseteq\mathcal{K}_{s}:=\mathcal{K}\cap\mathbb{E}_{s}$, and there exist scalars $s,s_{\mathcal{K}}^{*}\geq 1$ such that,
@@
    for all ${\mathbf{x}},{\mathbf{y}}\in\mathbb{E}_{s}$: $\|{{\mathbf{x}}-{\mathbf{y}}}\|\leq\sqrt{2s}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}$,
@@
    for all ${\mathbf{x}}\in\mathcal{K}$ and ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$: $\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|\leq\sqrt{s^{*}_{\mathcal{K}}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}$,
@@
    for all ${\mathbf{x}}\in\mathcal{K}$: $\arg\min\nolimits_{{\mathbf{y}}\in\mathbb{E}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|=\arg\min\nolimits_{{\mathbf{y}}\in\mathbb{E}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|_{2}$ (note the argmin need not be a singleton),
In the context of Problems (1) and (2) we should think of $\mathbb{E}_{s}$ as the restriction of the space $\mathbb{E}$ to vectors with a bounded level of sparsity (either in the sense of number of non-zero entries or matrix rank), and accordingly $\mathcal{K}_{s}$ is the restriction of $\mathcal{K}$ to vectors in $\mathbb{E}_{s}$. Additionally, while not part of Assumption 1, we shall implicitly assume that it is computationally efficient to compute a Euclidean projection onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$ (in case the projection is not a singleton, compute one element in the set of possible projections). Accordingly, we denote by $\Pi_{\mathbb{E}_{s}}[\cdot]$ and $\Pi_{\mathcal{K}_{s}}[\cdot]$ the Euclidean projection operation onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$, respectively. Note that due to Item 3 in Assumption 1, $\Pi_{\mathbb{E}_{s}}[\cdot]$ is also the projection w.r.t. the norm $\|{\cdot}\|$. In the context of both Problems (1) and (2), we shall demonstrate in the sequel that computing these projections is much more efficient than computing the Euclidean projection over the entire feasible set $\mathcal{K}$ when the optimal solutions are indeed sparse.
We can now present our algorithm for solving the generic Problem (5) (under the assumptions listed above). Our algorithm is composed of three simple ingredients. First, we use a Frank-Wolfe-like scheme [16] in which the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ are updated by taking a convex combination between the current feasible iterate and a new feasible point, as already shown in Eq. (4). Second, and differently from the classical Frank-Wolfe method in which the new point ${\mathbf{v}}_{t}$ is the output of a linear optimization oracle, here we take it to be a Euclidean proximal gradient step w.r.t. restricted set $\mathcal{K}_{s}$ (instead of w.r.t. the original set $\mathcal{K}$). This is similar to the approach taken in previous works [1, 11, 12, 13] (also referred to as a weak proximal oracle [11, 13]). Third, the $\ell_{2}$-term in this proximal gradient step is not w.r.t. the current iterate ${\mathbf{x}}_{t}$, but with respect to its projection onto $\mathbb{E}_{s}$, which should be understood as a hard-thresholding (or sparsification) of ${\mathbf{x}}_{t}$.
@@
  ${\mathbf{v}}_{t}\leftarrow\arg\min\nolimits_{{\mathbf{v}}\in\mathcal{K}_{s}}\langle{{\mathbf{v}}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+2s\beta\eta\|{{\mathbf{v}}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}$ {equivalently, ${\mathbf{v}}_{t}\leftarrow\Pi_{\mathcal{K}_{s}}[\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta}\nabla{}f({\mathbf{x}}_{t})]$}
  set $\gamma_{t}\in[0,1]$ either by line-search: $\gamma_{t}\leftarrow\arg\min_{\gamma\in[0,1]}f((1-\gamma){\mathbf{x}}_{t}+\gamma{\mathbf{v}}_{t})$ OR use a pre-defined $\gamma_{t}\in[0,1]$
  ${\mathbf{x}}_{t+1}\leftarrow(1-\gamma_{t}){\mathbf{x}}_{t}+\gamma_{t}{\mathbf{v}}_{t}$